OpenStudy (anonymous):

graph and solve 2x-5y=-5 ?

OpenStudy (kewlgeek555):

Okay, you have to find the x-value and then the y-value. Do you know how to do that?

OpenStudy (kewlgeek555):

Anyways. to find x, you simply input 0 for the value of y and solve. Same thing with finding y. When finding y, you input 0 into x and solve. So I'll help you with solving x.

OpenStudy (kewlgeek555):

\[2x-5y=-5\]\[2x-5 \times 0=-5\]\[2x - 0=-5\]\[-2x=-5\]2.5

OpenStudy (kewlgeek555):

So x = (2.5, 0)

OpenStudy (kewlgeek555):

Can you solve y?

OpenStudy (e.cociuba):